14. Choose the best Lewis structure for CH2Cl2. A) H--•-•-H H H-H-ċ-ċ-ċ:


H-H-C-CI-d: ca-C-C1 itl=3DCDCI: HICH
Answer #1

14. Choose the best Lewis structure for CH2Cl2. HL-0-0-H total valence electrons = 4 + 2 + (2 x 7) = 20 in chlorine there 3 u